# From Laboratory to Real World: A New Benchmark Towards Privacy-Preserved Visible-Infrared Person Re-Identification

### v1

@@ Abstract
Aiming to match pedestrian images captured under varying lighting conditions, visible-infrared person re-identification (VI-ReID) has drawn intensive research attention and achieved promising results. However, in real-world surveillance contexts, data is distributed across multiple devices/entities, raising privacy and ownership concerns that make existing centralized training impractical for VI-ReID. To tackle these challenges, we propose L2RW, a benchmark that brings VI-ReID closer to real-world applications. The rationale of L2RW is that integrating decentralized training into VI-ReID can address privacy concerns in scenarios with limited data-sharing regulation. Specifically, we design protocols and corresponding algorithms for different privacy sensitivity levels. In our new benchmark, we ensure the model training is done in the conditions that: 1) data from each camera remains completely isolated, or 2) different data entities (e.g., data controllers of a certain region) can selectively share the data. In this way, we simulate scenarios with strict privacy constraints which is closer to real-world conditions. Intensive experiments with various server-side federated algorithms are conducted, showing the feasibility of decentralized VI-ReID training. Notably, when evaluated in unseen domains (i.e., new data entities), our L2RW, trained with isolated data (privacy-preserved), achieves performance comparable to SOTAs trained with shared data (privacy-unrestricted). We hope this work offers a novel research entry for deploying VI-ReID that fits real-world scenarios and can benefit the community.
1 Introduction
The growing need for dependable pedestrian identification across all hours has made visible-infrared person re-identification (VI-ReID) a key technology [38, 49, 46]. It solves the challenge of identifying individuals in different spectral modalities, such as visible during the day and infrared at night. However, although existing methods [47, 48, 51, 30, 40, 41, 42, 39] achieve encouraging advancements, their deployment in real-world scenarios remains limited.
@@ 1 Introduction
Figure 1: Illustration of existing setting and the proposed L2RW setting. (a) The existing setting relies on centralized training, allowing unrestricted data sharing between cameras or entities, which brings serious privacy concerns. (b) Unlike this, our L2RW framework introduces camera-level and entity-level privacy regulation, simulating real-world privacy-preserved VI-ReID scenes.
In our view, the current experimental setting of the VI-ReID in the field is idealized and impedes its advancement to real-world applications. Specifically, current approaches [17, 3, 52, 25, 27, 31, 41, 40, 15] rely predominantly on centralized training, where data from multiple sources is aggregated on a central platform. While beneficial for learning modality-shared representations, they neglect practical constraints, especially in privacy and ownership concerns. In real-world scenarios, data is stored in a distributed manner, surveillance images are subject to strict privacy restrictions, etc. This centralized training not only complicates data sharing but also increases the risk of data leakage, bringing serious privacy concerns in real-world deployment.
Based on the above observation, we argue that it is essential to establish a benchmark that protects data privacy while still enabling effective VI-ReID. Therefore, we propose L2RW (Laboratory to Real World), a benchmark designed to simulate real-world VI-ReID scenarios with privacy preservation at its core. In L2RW, privacy is not an afterthought but a fundamental consideration. However, in reality, determining which data should be protected for privacy remains a challenging issue. For example, 1) pedestrian images captured by surveillance cameras are typically independent and without direct data sharing across cameras. 2) The entity, e.g., the data owner of a certain region, has the right to access all data. When different entities collaborate, should their data be isolated or kept shared?
This complexity brings us to a natural question in designing the L2RW benchmark for real-world VI-ReID: How can we define a proper and feasible privacy framework to adapt to different privacy levels? Thus, we design three protocols, i.e., Camera Independence (CI), Entity Independence (EI), and Entity Sharing (ES), for flexible privacy level control. For CI, all camera data remains isolated, simulating the highest privacy level where no data sharing occurs across every single camera as shown in the camera level of Fig. 1(b). Regarding a more relaxed privacy level in which all entity data is isolated while data can be shared inside each data entity, we propose the EI protocol as shown in the entity level of Fig. 1(b). When all data entities can be shared for centralized training, we name this protocol ES as shown in Fig. 1(a), which is the commonly used protocol in previous Re-ID methods. In the ES protocol, existing VI-ReID methods [44, 43, 26, 49, 14] can be seamlessly applied. However, note that these algorithms are designed for a closed-world setting, where training and testing are conducted on the same entity. To avoid this, in our ES protocol, we train on multiple shared entities but directly test on an unseen entity.
Evidently, existing VI-ReID methods cannot be directly applied under these privacy-preserving protocols (CI and EI) due to many technical obstacles: 1) In CI, a single camera may consist of only visible, only infrared, or both modalities, this modality mismatch between different cameras makes the training extremely challenging, leading to a modality incomplete issue. 2) In EI, since pedestrians vary across different entities, this non-sharing approach introduces the identity missing issue. 3) In both the CI and EI protocols, differences in data distribution across cameras or entities inevitably bring the domain shift issue.
To address the above challenges and make the L2RW benchmark actionable, we provide a novel algorithm customized for this task, named Decentralized Privacy-Preserved Training (DPPT) as a baseline solution. Inspired by the substantial success of federated learning [23, 19, 37, 18, 10, 13, 11], our DPPT employs decentralized training to ensure privacy protection. To solve the modality incomplete issue, DPPT converts the existing two-stream architecture into a one-stream architecture and modifies the current sampling way to eliminate the need for modality distinction. To handle the identity missing and domain shift issues, we design a memory rectification bank (MRB) that selects the top K features closest to the center point within each class as representatives. By averaging these top K features, we correct the center point to better reflect the current dominant class distribution. Then, MRB averages the dominant centers across clients and stores them in a global memory, which serves as a fair convergence point. This enables different clients to optimize towards the same goal. Generally, the contributions of this paper are summarized as follows:
  • We propose a benchmark named L2RW to simulate privacy-preserving real-world VI-ReID scenarios. Within the L2RW benchmark, we design three protocols, i.e., CI (Camera Independence), EI (Entity Independence) and ES (Entity Sharing). This simulates scenarios with different privacy constraints, bringing VI-ReID evaluation closer to real-world conditions.
  • We analyze the challenges encountered under the CI and EI protocols and propose a unified method named DPPT, which is the first work that handles privacy concerns for VI-ReID in a decentralized manner.
  • In our L2RW benchmark, unlike existing methods that validate solely on a single dataset, we merge several existing datasets and conduct the evaluation in a cross-domain manner to simulate the real-world scenarios.
  • Extensive experiments on three public VI-ReID datasets confirm the feasibility of decentralized training in L2RW, with our method achieving significant improvement on various federated learning baselines under CI. We also show that DPPT achieves performance under EI comparable to that of ES.
2 Related Work
2.1 Visible-Infrared Person ReID
@@
VI-ReID aims to match visible and infrared images of the same pedestrian captured by non-overlapped cameras. Many studies have emerged recently, achieving remarkable progress. For instance, Zhang et al. [49] explored diverse modality-shared features to mine significant cross-modality patterns. Jiang et al. [14] designed domain shifting (DNS) that augments modality-specific and modality-shared representation, thereby regulating the model to concentrate on consistency between modalities. Ren et al. proposed a novel implicit discriminative knowledge learning (IKDL) network to discover identity-aware salient information for aligning visible and infrared modalities. Some auxiliary-based methods [6, 48, 50] have also been developed to achieve modality alignment with the help of generated auxiliary modalities. Feng et al. [6] proposed a shape-guided diverse feature learning framework (SGIEL) that employs the body shape as the auxiliary information for modality alignment. Yu et al. [48] proposed a modality unifying network (MUN) that constructed a robust auxiliary modality by intra- and inter-modality learners.
@@
However, all the above methods are achieved under idealized lab settings, which neglect privacy constraints and heterogeneous environments pose significant challenges in reality. More importantly, these methods require paired visible and infrared images, making privacy-preserved implementation infeasible, especially under CI, where modality information is unknown. In addition, existing works neglect the research on unseen environments, struggling to generalize effectively in real-world scenes. Therefore, we propose L2RW, offering a novel practical and privacy-preserving benchmark for bringing VI-ReID closer to real world.
2.2 Federated Learning
@@
Federated Learning is a technology that enables multiple devices to collaborative learning without sharing private data. The pioneering work, FedAvg [23], averages the gradients of locally trained models and redistributes them to local clients for further training. Based on FedAvg, Li et al. [19] proposed FedProx, introducing a proximal term to ensure the convergence of the network in federated environments. Subsequently, Li et al. [18] designed MOON, which utilizes the modality representation similarity to correct the local training of clients. These methods brought new insights to the community, inspiring numerous impressive works that solve data heterogeneity [12, 28, 29, 34] and model heterogeneity [10, 22, 45] in federated learning. This progress provides solid technical support for our L2RW.
However, despite these advancements, existing federated learning methods cannot be directly applied to VI-ReID as they cannot handle specific issues such as modality incomplete, identity missing, and domain shift encountered in real-world scenarios.
2.3 Domain Generalization
@@
Domain generalization (DG) uses multiple seen domains to train a model that generalizes well to unseen domains. This technology aims to address the key issue of current deep learning methods, which heavily rely on the assumption that training and testing sets are independently and identically distributed. In practice, new information often arrives in continuous portions, leading to a domain gap between newly acquired data and the original data. This gap makes it challenging for the trained model to perform well on unseen data. Numerous works have emerged to extract domain-invariant features via contrastive loss [7, 9, 2], adversarial learning [16, 20, 4], causal learning [21, 1, 33], etc. Despite their success, these methods require centralized data from source domains and assume that data within the same domain share the same distribution. These not only limit the applicability of these methods in decentralized settings but also increase the risk of privacy leakage.
@@
Conversely, our designed EI protocol in the L2RW benchmark eliminates the need to share source domain data, addressing privacy concerns. This also shows the feasibility of privacy domain generalization in VI-ReID and shows its potential for secure and effective deployment in reality.
3 Methodology
@@
In this section, we describe the proposed L2RW benchmark in detail. In brief, Sec. 3.1 will present the overview of the L2RW. Then, we elaborate on the three protocols we designed, which are grounded in real-world scenarios: camera independence (CI) in Sec. 3.2.1, entity independence (EC) and entity sharing (ES) in Sec. 3.2.2. Finally, we introduce our proposed decentralized privacy-preserved training (DPPT) in Sec. 3.3.
@@
Figure 2: Protocols and proposed method. (a) are the protocols of camera independence (CI) and entity independence (EI) for privacy-preserved VI-ReID. Another our proposed protocol entity sharing (ES) is omitted in the figure, cause its training way is the same as that used in existing VI-ReID methods. (b) is the local training of our proposed decentralized privacy-preserved training (DPPT).
3.1 Overview
@@
The overall pipeline of the L2RW benchmark is shown in Fig. 2(a). Let $D_{m}=\{(x_{i},y_{i})\}_{i=1}^{N_{m}}$ denote the m-th local private data, where $x_{i}$ and $y_{i}$ denote the pedestrian images and corresponding label, and $N_{m}$ is the local data scale. The overall training set is denoted as $D=\{D_{1},D_{2},...,D_{n}\}$. In the realistic VI-ReID, each private dataset $D_{m}$ can represent the data collected by a single camera or the surveillance data from within a specific entity. The overall process of the L2RW can be roughly summarized into five steps. ① Local Training: Each private data is trained by a local model $\theta^{l}$ (client). ② Upload: Each client model weight is uploaded to the server. ③ Aggregation: The uploaded local model weights are aggregated to obtain the global model weight $\theta^{g}$. ④ Download: The global model weight $\theta^{g}$ is downloaded by the clients for next local training. By repeating step ① to ④ until convergence, execute ⑤ Test: test seen domains or unseen domains according to the protocol.
3.2 Benchmark Protocols

### v2

# L2RW+: A Comprehensive Benchmark Towards Privacy-Preserved Visible-Infrared Person Re-Identification

@@ Abstract
Visible-infrared person re-identification (VI-ReID) is a challenging task that aims to match pedestrian images captured under varying lighting conditions, which has drawn intensive research attention and achieved promising results. However, existing methods adopt the centralized training, ignoring the potential privacy concerns as the data is distributed across multiple devices or entities in reality. In this paper, we propose L2RW+, a benchmark that brings VI-ReID closer to real-world applications. The core rationale behind L2RW+ is that incorporating decentralized training into VI-ReID can address privacy concerns in scenarios with limited data-sharing constrains. Specifically, we design protocols and corresponding algorithms for different privacy sensitivity levels. In our new benchmark, we simulate the training under real-world data conditions that: 1) data from each camera is completely isolated, or 2) different data entities (e.g., data controllers of a certain region) can selectively share the data. In this way, we simulate scenarios with strict privacy restrictions, which is closer to real-world conditions. Comprehensive experiments show the feasibility and potential of decentralized VI-ReID training at both image and video levels. In particular, with increasing data scales, the performance gap between decentralized and centralized training decreases, especially in video-level VI-ReID. In unseen domains, decentralized training even achieves performance comparable to SOTA centralized methods. This work offers a novel research entry for deploying VI-ReID into real-world scenarios and can benefit the community. Code is available at: https://github.com/Joey623/L2RW.
Index Terms:
Privacy-preserved, Visible-infrared, person re-identification.
I Introduction
The growing need for reliable pedestrian identification throughout the day has made visible-infrared person re-identification (VI-ReID) a crucial technology [2, 3, 4]. It solves the challenge of identifying individuals under varying lighting conditions, such as visible during the daytime and infrared at night. Although existing methods [5, 6, 7, 8, 9, 10, 11, 12] have achieved encouraging advances, in our view, their settings are too idealized, thus restricting real-world deployments. To be specific, current VI-ReID methods adopt the centralized training, where data from multiple sources are aggregated on a central platform. Such settings overlook practical constraints such as privacy and data ownership, as surveillance data are typically distributed, sensitive, and subject to strict privacy restrictions, etc. This centralized training way in VI-ReID is overly idealized, and increases the risk of data leakage during the transmission of pedestrian images to the central server, bringing serious privacy concerns in real-world deployment.
@@ I Introduction
Figure 1: Illustration of the existing setting and the proposed L2RW+ setting. (a) The existing setting relies on centralized training, allowing unrestricted data sharing between cameras or entities, which brings serious privacy concerns. Img. and Vid. denote the image and video, respectively. (b) Unlike this, our L2RW+ framework introduces camera-level and entity-level privacy regulation, simulating real-world privacy-preserved VI-ReID scenes. Our L2RW+ considers both image and video data, offering a comprehensive benchmark for VI-ReID.
To address this issue, our previous work presented L2RW [1] in CVPR2025, a benchmark that protects data privacy while sitll enabling effective VI-ReID. L2RW defined three protocols, i.e., Camera Independence (CI), Entity Independence (EI) and Entity Sharing (ES), to adapt to different privacy levels, thereby achieving flexible privacy level control. The definition of three protocols is shown in Tab. I. For CI, all camera data is isolated, which simulates the strictest privacy level where no data sharing occurs across every single camera as shown in the camera level of Fig. 1(b). Regarding a more relaxed privacy level in which all entity data is isolated while data can be shared inside each data entity, we propose the EI protocol as shown in the entity level of Fig. 1(b). When all data entities can be shared for centralized training, we name this protocol ES, as shown in Fig. 1(a), which is the commonly used protocol in previous Re-ID methods. In the ES protocol, existing VI-ReID methods [13, 14, 15, 3, 16] can be seamlessly applied. However, these methods assume a closed-world setting where training and testing data come from the same entity, overlooking the domain generalization needed for open-world deployments. To address this, we evaluate generalization by training on multiple entities and testing on an unseen one under EI and ES protocols. Moreover, existing VI-ReID methods cannot be directly applied under these privacy-preserving protocols (CI and EI) due to modality incomplete, identity missing, and domain shift issues, as detailed in Sec. II-B. To solve this, our L2RW proposed the Image-based Decentralized Privacy-Preserved Training (I-DPPT), which employed decentralized training to ensure privacy protection while achieving VI-ReID. Generally, the contribution of L2RW can be summarized as follows: (1) We designed three protocols, i.e., CI (Camera Independence), EI (Entity Independence), and ES (Entity Sharing) to simulate scenarios with different privacy constraints; (2) We identified challenges under the CI and EI protocols, and proposed I-DPPT, the first decentralized VI-ReID method addressing privacy concerns; (3) Unlike existing methods evaluated on single datasets, our L2RW benchmark is the first work to merge multiple datasets for cross-domain evaluation. (4) Extensive experiments on three publicly available image-based VI-ReID datasets confirmed the feasibility of decentralized training in L2RW, with our method achieving significant improvement on various federated learning baselines under CI.
TABLE I: The details of three protocols in L2RW benchmark.
[…]
EI Data is shared within an entity but isolated across entities Medium
ES All entity data is shared Low
Despite these contributions, we argue that L2RW remains limited in two key aspects. First, L2RW only supports image data, whereas real-world surveillance systems predominantly collect continuous video streams rich in temporal and motion information, which cannot be processed by L2RW. Second, L2RW only demonstrates the feasibility and generalization ability of image-based VI-ReID, but lacks evaluation under real-world conditions such as environmental noise, adversarial attacks, and privacy leakage risks, leaving the model’s robustness and security largely unexplored.
To address these challenges, we propose L2RW+, an extension of L2RW that introduces the following improvements:
  • From a methodology perspective, we extend decentralized VI-ReID from image-level to video-level and design Video-based Decentralized Privacy-Preserved Training (V-DPPT) for video data. V-DPPT effectively leverages temporal cues while respecting privacy constraints, enabling accurate retrieval from video data. L2RW+ unifies image and video evaluation under all three privacy protocols (CI, EI, ES). To the best of our knowledge, this is the first work that offers the comprehensive benchmark for privacy-preserved VI-ReID.
  • From experiments, we demonstrate the feasibility and potential of decentralized training in VI-ReID. Specifically, under the CI protocol, our proposed DPPT shows the retrieval ability, even surpassing some methods trained under the centralized manner. Notably, the performance gap between decentralized and centralized training narrows on video data, suggesting that temporal information is useful for decentralized setups. Furthermore, under the EI and ES protocols, the model trained in a decentralized manner exhibits stronger generalization to unseen environments, especially when more training data becomes available. These findings not only validate the practical viability of decentralized learning in privacy-preserving VI-ReID but also emphasize the need for larger and more diverse VI-ReID datasets to better understand the consistent representation of visible and infrared modalities. Our L2RW+ offers meaningful insights and establishes a solid foundation for future privacy-preserved VI-ReID research.
  • From a practical perspective, we show through extensive experiments that decentralized training is more robust than centralized training. In real-world VI-ReID applications, systems often face challenging conditions like rain, snow, or fog, and must also handle sensitive identity data that remains vulnerable to attacks. To reflect these challenges, we add noise to simulate harsh environments and apply various adversarial attacks. The results show that decentralized training maintains more stable performance and offers better privacy protection, as confirmed by gradient inversion experiments where pedestrian attributes are harder to recover.
Generally, this is the first work to introduce decentralized training into VI-ReID across both image and video data, showing superior performance on privacy, generalization, and robustness over centralized approaches. We believe our findings offer a practical and promising direction for deploying VI-ReID in real-world surveillance systems.
II Related Work
II-A Visible-Infrared Person ReID
VI-ReID aims to match visible and infrared images of the same pedestrian captured by non-overlapping cameras. Many studies have emerged recently, achieving remarkable progress. Based on the input type, VI-ReID can be categorized into image-based and video-based subtasks.
@@ II-A Visible-Infrared Person ReID
Image-based VI-ReID. Zhang et al. [3] explored diverse modality-shared features to mine significant cross-modality patterns. Jiang et al. [16] designed domain shifting (DNS) that augments modality-specific and modality-shared representation, thereby regulating the model to concentrate on the consistency between modalities. Ren et al. proposed a novel implicit discriminative knowledge learning (IKDL) network to discover identity-aware salient information for aligning visible and infrared modalities. Some auxiliary-based methods [17, 6, 18] have also been developed to achieve modality alignment with the help of generated auxiliary modalities. Feng et al. [17] proposed a shape-guided diverse feature learning framework (SGIEL) that employs the body shape as the auxiliary information for modality alignment. Yu et al. [6] proposed a modality unifying network (MUN) that constructed a robust auxiliary modality by intra- and inter-modality learners.
Video-based VI-ReID. Video-based VI-ReID aims to match pedestrian sequences across visible and infrared modalities. Different from image-based VI-ReID, it contains extra temporal information. Liu et al. [19] built the first video VI-ReID dataset dubbed HITSZ-VCM and designed the MITML method, attracting increasing research interest of this task. Specifically, Li et al. [20] designed an Intermediary-guided Bidirectional spatial–temporal Aggregation Network (IBAN) to mitigate the modality gap and learn the spatial-temporal representation by leveraging anaglyph data as the auxiliary. Du et al. [21] proposed AuxNet, which employs GANs for modality alignment and adopts a curriculum learning strategy to learn temporal information from the sequence, effectively reducing the cross-modality gap. Yu et al. [22] introduced TF-CLIP, which extracts identity-specific sequence features to replace the original text embeddings used in CLIP. Li et al. [23] introduced VLD that learned modality-invariant sequence-level pedestrian representations guided by video-level language prompts.
@@
However, all the above methods are achieved under idealized lab settings, which neglect privacy constraints and heterogeneous environments that pose significant challenges in reality. More importantly, these methods require paired visible and infrared images, making privacy-preserved implementation infeasible, especially under CI, where modality information is unknown. In addition, existing works neglect the research on unseen environments, struggling to generalize effectively in real-world scenes. Therefore, we propose L2RW+, offering a novel, practical, and privacy-preserving benchmark for pushing VI-ReID closer to real world settings.
II-B Federated Learning
@@
Federated Learning (FL) is a technology that enables multiple devices to collaborative learning without sharing private data. The pioneering work, FedAvg [24], averages the gradients of locally trained models and redistributes them to local clients for further training. Based on FedAvg, Li et al. [25] proposed FedProx, introducing a proximal term to ensure the convergence of the network in federated environments. Subsequently, Li et al. [26] designed MOON, which utilizes the modality representation similarity to correct the local training of clients. These methods brought new insights to the community, inspiring numerous impressive works that solve data heterogeneity [27, 28, 29, 30] and model heterogeneity [31, 32, 33] in federated learning. This progress provides solid technical support for our L2RW+.
However, despite these advancements, existing FL methods are general methods and ignore the unique challenges arising from the heterogeneous and privacy-sensitive nature of VI-ReID data, as illustrated in Fig. 2. Let $x$ and $y$ denote the pedestrian images and their corresponding identity labels. Each client refers to an individual camera or data entity participating in decentralized training without data sharing. The main challenges include:
  • Domain shift: $P_{i}(x|y)\neq P_{j}(x|y)$. Pedestrian images on different clients exhibit the non-iid (Independent and identically distributed) issue, where the corresponding feature distributions are different across different clients.
  • Identity missing: $P_{i}(y)\neq P_{j}(y)$. The same subject might not appear in all cameras or entities.
To address these challenges, we propose two decentralized training frameworks: I-DPPT for image-based VI-ReID and V-DPPT for video-based VI-ReID. The DPPT is designed to work under privacy constraints and directly handle the problems of modality incompleteness, identity missing, and domain shift in federated settings. In addition, by adapting the model structure and training strategy to the characteristics of VI-ReID data, our approach achieves effective person retrieval without sharing data across clients.
Figure 2: Illustration of specific challenges in privacy-preserved VI-ReID. Each single camera may consist of only visible, only infrared, or both modalities, we name this modality incomplete. The pedestrian may not captured by all cameras or entities, which we define as identity missing. Data distribution is different across cameras/entities, inevitably resulting in domain shift.
II-C Domain Generalization
@@
Domain generalization (DG) uses multiple seen domains to train a model that generalizes well to unseen domains. This technology aims to address the key issue of current deep learning methods, which heavily rely on the assumption that training and testing sets are independently and identically distributed. In practice, new information often arrives in continuous portions, leading to a domain gap between newly acquired data and the original data. This gap makes it challenging for the trained model to perform well on unseen data. Numerous works have emerged to extract domain-invariant features via contrastive loss [34, 35, 36], adversarial learning [37, 38, 39], causal learning [40, 41, 42], etc. Despite their success, these methods require centralized data from source domains and assume that data within the same domain shares the same distribution. These not only limit the applicability of these methods in decentralized settings but also increase the risk of privacy leakage.
@@
Conversely, our designed EI protocol in the L2RW+ benchmark eliminates the need to share source domain data, addressing privacy concerns. This also shows the feasibility of privacy domain generalization in VI-ReID and shows its potential for secure and effective deployment in reality.
III Methodology
@@
In this section, we describe the proposed L2RW+ benchmark in detail. Sec. III-A presents the overview of the L2RW+. Then, three protocols we designed are elaborated in Sec. III-B, which are grounded in real-world scenarios: camera independence (CI), entity independence (EI), and entity sharing (ES). Finally, we introduce our proposed decentralized privacy-preserved training (DPPT) for image-based and video-based VI-ReID in Sec. III-C and Sec. III-D, respectively.
@@
Figure 3: Protocols and proposed method. (a) The protocols of camera independence (CI) and entity independence (EI) for privacy-preserved VI-ReID. Our proposed another protocol entity sharing (ES) is omitted in the figure, because its training way is the same as that used in existing VI-ReID methods. (b) The local training of our proposed image-baesd decentralized privacy-preserved training (I-DPPT). The local training of V-DPPT is shown in Fig. 4.
III-A Overview
@@
The overall pipeline of the L2RW+ benchmark is shown in Fig. 3(a). Let $D_{m}=\{(x_{i},y_{i})\}_{i=1}^{N_{m}}$ denote the m-th local image-level private data, where $x_{i}$ and $y_{i}$ denote the pedestrian images and corresponding label, and $N_{m}$ is the local data scale. The overall training set is denoted as $D=\{D_{1},D_{2},...,D_{n}\}$. In the realistic VI-ReID, each private dataset $D_{m}$ can represent the data collected by a single camera or the surveillance data from within a specific entity. The overall process of the L2RW+ can be roughly summarized into five steps. ① Local Training: Each private data is trained by a local model $\theta^{l}$ (client). ② Upload: Each client model weight is uploaded to the server. ③ Aggregation: The uploaded local model weights are aggregated to obtain the global model weight $\theta^{g}$. ④ Download: The global model weight $\theta^{g}$ is downloaded by the clients for next local training. By repeating step ① to ④ until convergence, execute ⑤ Test: test seen domains or unseen domains according to the protocol.
III-B Benchmark Protocols